# EUROPEAN ORGANIZATION FOR NUCLEAR RESEARCH (CERN) CERN-EP-2025-104 LHCb-PAPER-2025-007

### v2

@@
A study on the Bose–Einstein correlations for triplets of same-sign pions is presented. The analysis is performed using proton-proton collisions at a centre-of-mass energy of = 7 TeV, recorded by the LHCb experiment, corresponding to an integrated luminosity of 1.0. For the first time, the results are interpreted in the core-halo model. The parameters of the model are determined in regions of charged-particle multiplicity. This measurement provides insight into the nature of hadronisation in terms of coherence, showing a coherent emission of pions.
Submitted to JHEP
@@ 1 Introduction
Previous results on three-particle correlations have already been reported by the ALICE experiment at the LHC [10]. However, with a limited interpretation of the results within the core-halo model [9]. Recently, measurements of three-particle correlations were interpreted within the core-halo model (see Sec. 2.4) at the PHENIX experiment [11, 12]. This model provides a set of parameters describing the expansion of the system after a collision and describes hadronic source properties not available in two-particle correlation analyses.
This paper reports the measurements of the three-particle correlation strength [11, 13] in different regions of charged-particle multiplicity, employing collisions at = 7 TeV recorded by the LHCb experiment, corresponding to an integrated luminosity of 1 fb-1. A convolution of the information from the two-particle correlation function, determined in a previous LHCb analysis [14], with the three-particle BEC, may probe the parameters of the core-halo model [9], disentangling the partial coherence and the halo contribution. In particular, the parameters can be studied in relation to a thermalised core. To achieve this, a new parameter is introduced, [11], being a function of both two- and three-particle correlation strengths, and . This new parameter is different from unity when additional effects are present within the core, such as an imperfectly thermalised core or partial coherence. The analysis investigates whether reflects such additional effects. The values of the partial coherence parameter and the fraction of the particles that originate from the core are also measured as a function of the charged-particle multiplicity.
2 Analysis method
@@ 3 Detector and dataset
The LHCb detector [26, 27] is a single-arm forward spectrometer designed to study particles containing or quarks. It features a high-precision tracking system, which includes a silicon-strip vertex detector (VELO) [28] surrounding the interaction region, covering the pseudorapidity range . Upstream of a dipole magnet with a bending power of about 4 T m, a large-area silicon-strip detector is located, along with three stations of silicon-strip detectors and straw drift tubes [29] downstream of the magnet. The tracking system measures the momentum, , with a relative uncertainty ranging from 0.5% at low momentum to 1% at 200 GeV. The impact parameter (IP), or the minimum distance of a track to a primary vertex (PV), is measured with a resolution of (15 + 29/), where is the transverse component of in GeV. Charged hadrons are identified using information from two ring-imaging Cherenkov detectors [30]. Photons, electrons and hadrons are identified by a calorimeter system. Muons are detected by a system of alternating layers of iron and multiwire proportional chambers [31]. The trigger [32] consists of a hardware stage based on information from the calorimeter and muon systems, followed by a software stage that performs full event reconstruction.
In this analysis, a collision dataset of no-bias and minimum-bias triggered events is used, collected in 2011 at a centre-of-mass energy of = 7 TeV, corresponding to an integrated luminosity of 1.0. The no-bias trigger selects events randomly, while the minimum-bias trigger requires at least one reconstructed VELO track.222A VELO track is reconstructed only with hits registered in the VELO detector. The data were collected with an average number of visible interactions per bunch crossing (pile-up) of 1.4 [33]. To avoid biases related to trigger requirements, a sample of uncorrelated interactions is constructed as described in Sec. 3.
In the simulation, collisions are generated using Pythia 8 [34] with a specific LHCb configuration [35], excluding the BEC effect. The decays of hadronic particles are simulated using EvtGen [36], with final-state radiation generated by Photos [37]. The interaction of the generated particles with the detector and its response are modelled using the Geant4 toolkit [38, 39], as described in [40]. For the study of systematic effects, an additional sample is simulated using Pythia 6.4 [41] with the Perugia0 tune [42].
@@ 4 Event selection
The following single-particle selection criteria are applied. All pion candidates must have reconstructed track segments in the VELO, with , and in tracking stations downstream of the magnet. Each track is required to have a good-quality fit [43], a transverse momentum GeV, and no associated signal in the muon stations. Both pion candidates must be assigned to the same PV. Particles are assigned to the PV that minimises the value of the impact parameter, defined as the difference in the vertex-fit values of a given PV reconstructed with and without the track in question. A loose requirement on the track impact parameter (IP 0.4 mm) is applied to retain the majority of the particles originating from the selected PV. To reduce contamination from fake and clone tracks,333Fake tracks are those which do not correspond to any particle trajectory, but are reconstructed from a number of unrelated hits. Clone tracks are multiple tracks reconstructed from hits which were deposited by a single charged particle. only the track with the best is retained when multiple tracks share the same hits in the VELO subdetector. Additionally, fake tracks are removed by applying requirements on the track and using the output of a dedicated neural network [30].
@@
Particle identification (PID) is performed using a neural network that incorporates subdetector information to calculate the probability for a particle to be identified as a specific type [30]. The simulated PID quantities are corrected using calibration samples from the data [44]. Corrections of PID variables among particles are also taken into account. It is crucial to maintain a sample with high purity; however, applying a strict requirement on the probability for a particle to be identified as a pion, , could suppress low-momentum pions that contribute to the BEC effect and significantly impact the signal region of the correlation function. To ensure consistency with previous two-pion analyses of collisions [14], the requirement of 0.65 is applied. Varying this threshold from 0.5 to 0.8 does not result in significant changes in the measured correlation function.
@@ 5 Fitting the correlation function
Finally, the distributions of the double ratio of correlation functions for same-sign pion triplets, , using the event-mixed reference sample (as described in Sec. 2.1), are fitted in three different regions (bins) of VELO track multiplicity per collision using the parameterisation in Eq. 13 with four-momenta differences , , selected to fall into the same -bin. A binned maximum-likelihood fit is used. The fitted parameters are , , and , while the values of and are taken from the published two-pion BEC analysis [14], whose values are, depending of bin, in the range 1.01–1.80 fm and 0.57–0.72, respectively. The results of the fits to the three-particle double ratios are presented in Fig. 1. The Bose–Einstein correlation effect is evident as a signal enhancement for low values of . It is worth noting that the fit quality in BEC studies is not expected to be perfect. Due to the lack of theoretical description of nonfemtoscopic background contributions as well as the compromise between the fit quality and interpretability of the measured correlation parameters, the obtained values of per degree of freedom are often larger than unity.
@@ 7 Results
The analysis most closely aligned in scope with the present study was performed by the PHENIX experiment at RHIC [11, 12]. It uses the core-halo model to interpret BEC for triplets of pions in Au-Au collisions at in bins of transverse mass,444Transverse mass , where is the average pair transverse momentum. measuring and correlation strength . Despite the differences in the experimental setups, the central values of display a similar tendency to the results of the present study, which could indicate a partially coherent emission. However, these central values are consistent with unity within statistical and systematic uncertainties, preventing definitive conclusions. The reported values of three-particle correlation strengths are higher than those measured for collisions in the present analysis. The PHENIX analysis does not present the dependence on particle multiplicity, therefore, a direct comparison of the results for and Au-Au is not possible.
@@ 8 Conclusions
The correlations among three indistinguishable pions are studied using data collected by the LHCb experiment in 2011 at a centre-of-mass energy of 7 TeV in proton-proton collisions. The double-ratio technique is employed to correct for nonfemtoscopic background, together with the Gamov penetration factor utilised to account for electromagnetic final-state interactions that can be factorised using the generalised Riverside method. The fits applied to the three-pion double ratios of the correlation function provide parameters used to interpret the results within the frame of the core-halo model. The values of the core-halo parameters are studied in three activity classes related to the charged-particle multiplicity. The measured dependencies of the core-halo parameters on the charged-particle multiplicity suggest a partially coherent emission of pions. In particular, the partial coherence parameter increases with multiplicity, indicating coherent emission of pions. At the same time, the parameter, which is systematically below unity for higher-multiplicity bins, further suggests a partially coherent emission. However, due to the large uncertainties, it is also consistent with unity, preventing any firm conclusions from being drawn in this case.

### v4

@@
A study on the Bose–Einstein correlations for triplets of same-sign pions is presented. The analysis is performed using proton-proton collisions at a centre-of-mass energy of = 7 TeV, recorded by the LHCb experiment, corresponding to an integrated luminosity of 1.0. For the first time, the results are interpreted in the core-halo model. The parameters of the model are determined in regions of charged-particle multiplicity. This measurement provides insight into the nature of hadronisation in terms of coherence, being consistent with the presence of coherent emission of pions.
Published in JHEP 08 (2025) 174
@@ 1 Introduction
Previous results on three-particle correlations have already been reported by the ALICE experiment in PbPb collisions at the LHC [10]. However, with a limited interpretation of the results within the core-halo model [9]. Recently, measurements of three-particle correlations in AuAu collisions were interpreted within the core-halo model (see Sec. 2.4) at the PHENIX experiment [11, 12]. This model provides a set of parameters describing the expansion of the system after a collision and describes hadronic source properties not available in two-particle correlation analyses.
This paper reports for the first time the measurements of the three-particle correlations in small collision systems, performed in different regions of charged-particle multiplicity, employing collisions at = 7 TeV recorded by the LHCb experiment, corresponding to an integrated luminosity of 1 fb-1. A convolution of the information from the two-particle correlation function, determined in a previous LHCb analysis [13], with the three-particle BEC, may probe the parameters of the core-halo model [9], disentangling the partial coherence and the halo contribution. In particular, the parameters can be studied in relation to a thermalised core. To achieve this, a new parameter is introduced, [11], being a function of both two- and three-particle correlation strengths, and [11, 14]. This new parameter is different from unity when additional effects are present within the core, such as an imperfectly thermalised core or partial coherence. The analysis investigates whether reflects such additional effects. The values of the partial coherence parameter and the fraction of the particles that originate from the core are also measured as a function of the charged-particle multiplicity. A key question is whether coherent emission can also be observed in small collision systems like , which would manifest in the multiplicity-dependence of and parameters.
The paper is organised as follows. In Sec. 2 an overview of the theory regarding the two- and three-particle correlations is given, followed by an introduction of the core-halo model used in the present analysis. The LHCb detector and data samples used are described in Sec. 3, while the event selection and the fits to the three-pion correlation functions are reported in Sec. 4 and 5, respectively. The systematic uncertainties and final results are presented in Secs. 6 and 7. A summary of the studies performed is given in Sec. 8.
2 Theoretical framework
@@ 3 Detector and dataset
The LHCb detector [26, 27] is a single-arm forward spectrometer designed to study particles containing or quarks. It features a high-precision tracking system, which includes a silicon-strip vertex detector (VELO) [28] surrounding the interaction region, covering the pseudorapidity range .222Pseudorapidity is the variable that corresponds to the polar angle of the track with respect to the beam axis, . References to pseudorapidity in the present study correspond to the values in the LHCb laboratory frame. The LHCb coordinate system is right-handed, with the -axis pointing along the beam axis, -axis in the vertical direction, and -axis in the horizontal direction. The ()-plane is the bending plane of the dipole magnet. Upstream of a dipole magnet with a bending power of about 4 T m, a large-area silicon-strip detector is located, along with three stations of silicon-strip detectors and straw drift tubes [29] downstream of the magnet. The tracking system measures the momentum, , with a relative uncertainty ranging from 0.5% at low momentum to 1% at 200 GeV. The impact parameter, or the minimum distance of a track to a primary vertex (PV), is measured with a resolution of (15 + 29/), where is the transverse component of in GeV. Charged hadrons are identified using information from two ring-imaging Cherenkov detectors [30]. Photons, electrons and hadrons are identified by a calorimeter system. Muons are detected by a system of alternating layers of iron and multiwire proportional chambers [31]. The trigger [32] consists of a hardware stage based on information from the calorimeter and muon systems, followed by a software stage that performs full event reconstruction.
In this analysis, a collision dataset of no-bias and minimum-bias triggered events is used, collected in 2011 at a centre-of-mass energy of = 7 TeV, corresponding to an integrated luminosity of 1.0. The data sample available in the present study corresponds to events. The no-bias trigger selects events randomly, while the minimum-bias trigger requires at least one reconstructed VELO track.333A VELO track is reconstructed only with hits registered in the VELO detector. The data were collected with an average number of visible interactions per bunch crossing (pile-up) of 1.4 [33]. To avoid biases related to trigger requirements, a sample of uncorrelated interactions is constructed as described in Sec. 3.
In the simulation, collisions are generated using Pythia 8 [34] with a specific LHCb configuration [35], excluding the BEC effect. The decays of hadronic particles are simulated using EvtGen [36], with final-state radiation generated by Photos [37]. The interaction of the generated particles with the detector and its response are modelled using the Geant4 toolkit [38, 39], as described in [40].444The version numbers of key simulation tools: PYTHIA 8.1, EvtGen 2.2, PHOTOS 2.0, Geant4 v70r0p1. The simulated dataset contains minimum bias events. For the study of systematic effects, an additional sample of minimum bias events is simulated using Pythia 6.4 [41] with the Perugia0 tune [42].
In order to be compatible with the two-pion correlation analysis results [13], the same data and simulation samples, together with the same packages versions, are used.
@@ 4 Event selection
The following single-particle selection criteria are applied. All pion candidates must have reconstructed track segments in the VELO, with , and in tracking stations downstream of the magnet. Each track is required to have a good-quality fit [43], a transverse momentum GeV, and no associated signal in the muon stations. Both pion candidates must be assigned to the same PV. Particles are assigned to the PV that minimises the value of the impact parameter, defined as the difference in the vertex-fit values of a given PV reconstructed with and without the track in question. A loose requirement on the track impact parameter ( 0.4 mm) is applied to retain the majority of the particles originating from the selected PV. To reduce contamination from fake and clone tracks,555Fake tracks are those which do not correspond to any particle trajectory, but are reconstructed from a number of unrelated hits. Clone tracks are multiple tracks reconstructed from hits which were deposited by a single charged particle. only the track with the best is retained when multiple tracks share the same hits in the VELO subdetector. Additionally, fake tracks are removed by applying requirements on the track and using the output of a dedicated neural network [30].
@@
Particle identification (PID) is performed using a neural network that incorporates subdetector information to calculate the probability for a particle to be identified as a specific type [30]. The simulated PID quantities are corrected using calibration samples from the data [44]. Corrections of PID variables among particles are also taken into account. It is crucial to maintain a sample with high purity; however, applying a strict requirement on the probability for a particle to be identified as a pion, , could suppress low-momentum pions that contribute to the BEC effect and significantly impact the signal region of the correlation function. To ensure consistency with previous two-pion analyses of collisions [13], the requirement of 0.65 is applied. With this relatively loose selection criterion, the single-track pion identification purity is approximately 98%. This leads to a pion pair purity of 96%. The contamination from mixed pairs containing only one true pion is about 4%. This level of contamination has been verified to have no impact on the correlation function, as the Bose–Einstein correlation effect does not occur in such cases. The total contribution from same-type particle pairs not containing pions is below 10-4. Varying threshold from 0.5 to 0.8 does not result in significant changes in the measured correlation function.
@@ 5 Fitting the correlation function
Finally, the distributions of the double ratio of correlation functions for same-sign pion triplets, , using the event-mixed reference sample (as described in Sec. 2.1), are fitted. The fit is performed in three different regions (bins) of VELO track multiplicity per collision using the parameterisation in Eq. 13 with four-momenta differences , , selected to fall into the same -bin. A binned maximum-likelihood fit is used. The fitted parameters are , , and , while the values of and are taken from the published two-pion BEC analysis [13], whose values are, depending of bin, in the range 1.01–1.80 fm and 0.57–0.72, respectively. The results of the fits to the three-particle double ratios are presented in Fig. 1, while the related can be found in Appendix A. The Bose–Einstein correlation effect is evident as a signal enhancement for low values of . It is worth noting that the fit quality in BEC studies is not expected to be perfect. Due to the lack of theoretical description of nonfemtoscopic background contributions as well as the compromise between the fit quality and interpretability of the measured correlation parameters, the obtained values of per degree of freedom are often larger than unity.
@@ 7 Results
The analysis most closely aligned in scope with the present study was performed by the PHENIX experiment at RHIC [11, 12]. It uses the core-halo model to interpret BEC for triplets of pions in Au-Au collisions at in bins of transverse mass,666Transverse mass , where is the average pair transverse momentum. measuring and correlation strength . Despite the differences in the experimental setups, the central values of display a similar tendency to the results of the present study, which could indicate a partially coherent emission. However, these central values are consistent with unity within statistical and systematic uncertainties, preventing definitive conclusions. The reported values of three-particle correlation strengths are higher than those measured for collisions in the present analysis. The PHENIX analysis does not present the dependence on particle multiplicity, therefore, a direct comparison of the results for and Au-Au is not possible.
@@ 8 Conclusions
The correlations among three indistinguishable pions are studied using data collected by the LHCb experiment in 2011 at a centre-of-mass energy of 7 TeV in proton-proton collisions. The double-ratio technique is employed to correct for nonfemtoscopic background, together with the Gamov penetration factor utilised to account for electromagnetic final-state interactions that can be factorised using the generalised Riverside method. The fits applied to the three-pion double ratios of the correlation function provide parameters used to interpret the results within the frame of the core-halo model. The values of the core-halo parameters are studied in three activity classes related to the charged-particle multiplicity. The measured dependencies of the core-halo parameters on the charged-particle multiplicity suggest a partially coherent emission of pions. In particular, the partial coherence parameter increases with multiplicity, indicating coherent emission of pions. At the same time, the parameter, which is systematically below unity for higher-multiplicity bins, further suggests a partially coherent emission. However, due to the large uncertainties, it is also consistent with unity, preventing any firm conclusions from being drawn in this case. The behaviour of indicates that the fraction of particles that originated from the core of the emission volume does not depend strongly on the number of charged particles produced in the event.